OpenStudy (anonymous):

PLEASE HELPPP! solve for x in this equation 2x-1/6 = 4x+1/12 + 2x+2/2x+1

OpenStudy (anonymous):

x = 0

jhonyy9 (jhonyy9):

2x-1 4x+1 2x+2 2(2x+1)(2x-1) (4x+1)(2x+1) + 12(2x+2) ---- = ----- + ------ ------ ------------- = ------------------------ 6 12 2x+1 12(2x+1) 12(2x+1) 2(2x+1)(2x-1)=(4x+1)(2x+1)+12(2x+2) 2(4x^2 -1)=8x^2 +6x +1 +24x +24 8x^2 -2 =8x^2 +30x +25 -27 =30x /3 -9=10x -9 x= ---- = -0,9 10